OpenStudy (anonymous):

(x+1)(x+4)=14 is this correct: x - 1 = 14 or x + 4 = 14 +1 +1 - 4 - 4 x = 15 x = 10

OpenStudy (anonymous):

I got, x=13 x=10

OpenStudy (anonymous):

you do the opposite of what is in parenthesis

OpenStudy (anonymous):

In the first parenthesis you have (x+1) but transfered it x-1 when solving. Is this correct?

OpenStudy (anonymous):

(x+1)(x+4)=14 x+1=14, x+4+14 -1 -1 -4 -4 x=13 x=10

OpenStudy (anonymous):

so mistake in adding 1 instead of subracting. thx!

OpenStudy (anonymous):

Unfortunately that is not correct at all =/ it is not true that:\[(x+1)(x+4)=14\Longrightarrow x+1=14,x+4=14\]You will have to multiply out the (x+1)(x+4) first, then subtract 14 from both sides, then solve the problem.

myininaya (myininaya):

\[x^2+x+4x+4=14 => x^2+5x+4-14=0\]

myininaya (myininaya):

\[x^2+5x-10=0\]